A study investigating the relationship between age and annual medical expenses randomly samples 100 individuals in a city. It is hoped that the sample will have a similar mean age as the entire population. Complete parts a and b below.

a. If the standard deviation of the ages of all individuals in the city is σ=14​, find the probability that the mean age of the individuals sampled is within two years of the mean age for all individuals in the city.​ (Hint: Find the sampling distribution of the sample mean age and use the central limit theorem. You​ don't have to know the population mean to answer​ this, but if it makes it​ easier, use a value such as μ=35​.)

The probability is ____ . ​(Round to three decimal places as​ needed.)

b. Would the probability be​ larger, or​ smaller, if σ=5​?​Why?

If σ were smaller​, the standard deviation of the sampling distribution would be ____ (larger, smaller) , so the probability would be ____( smaller. larger).
Answer #1

solution: o=14 naloo a) p118-M) <2) - 2 14 | Too LZL 14} voo) Z = X-ll airns = P(-1.43 <Zc1.43) P1201.43) - P126-1.43) 0.923